OpenStudy (anonymous):

I am having trouble with completing a square, please walk me through for this problem... 2x^2+5x-2=0

hero (hero):

I believe he said to "walk him through" the steps, not post ALL of the steps.

OpenStudy (anonymous):

My apologies. I'll... fix it.

OpenStudy (campbell_st):

start by dividing every term by 2 x^2 + 5/2 x - 1 = 0 add 1 to both sides x^2 + 5/2x = 1 for a quadratic x^2 + bx + c then you need to calculate (b/2)^2 this will be (5/2/2)^2 = (5/4)^2 this term is then added to both sides of the equation x^2 + 5/2 x + (5/4)^2 = 1 + (5/4)^2 left hand side is a perfect square (x + 5/4)^2 = 1 + 25/16 (x + 5/4)^2 = 41/16

OpenStudy (anonymous):

You would like to have $22,000 in 4 years by making regular deposits at the end of every three months in an annuity that pays 3.75% compounded quarterly. How much of the $22,000 comes from interest? Round your answer to the nearest dollar.